# Smooth Tests for Normality in ANOVA

### v3

@@ 5 Simulation
The results of Experiment II are quite similar; see Tables 5, 6, and 7, as well as Figure 2. However, from Table 5, under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to exhibit power because in this case, $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi(Y_{ij}^{(1,2)}-8))]=0$. This illustrates that $N\hat{\Psi}_{K}^{2}$ may be inconsistent against certain alternatives, especially when $K$ is small. Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, $\{\hat{K}=1\}$ does not occur, consistent with Theorem 10 (3), which ensures that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H_{1}^{\prime}$ with $K_{0}=2$.
Table 2: Empirical rejection rates under $H_{0}$ in Experiment I, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
10 0.050 0.048 0.048 0.044 0.040 0.056 0.034
20 0.046 0.044 0.034 0.040 0.048 0.066 0.046
30 0.042 0.036 0.044 0.068 0.068 0.060 0.046
40 0.058 0.040 0.036 0.058 0.052 0.058 0.050
50 0.064 0.044 0.054 0.054 0.062 0.068 0.060
60 0.054 0.034 0.038 0.044 0.050 0.070 0.062
70 0.056 0.038 0.048 0.046 0.052 0.056 0.048
80 0.038 0.050 0.072 0.042 0.048 0.058 0.048
90 0.062 0.066 0.052 0.064 0.058 0.048 0.040
100 0.034 0.050 0.050 0.060 0.054 0.052 0.038
110 0.054 0.032 0.054 0.060 0.050 0.038 0.036
120 0.052 0.052 0.038 0.044 0.052 0.052 0.050
130 0.052 0.050 0.064 0.060 0.058 0.068 0.064
140 0.062 0.040 0.026 0.046 0.030 0.038 0.030
150 0.054 0.044 0.062 0.034 0.058 0.056 0.048
Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment I
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.982 0.018 0 0 0
20 0.988 0.010 0 0 0.002
30 0.986 0.012 0 0.002 0
40 0.984 0.014 0.002 0 0
50 0.988 0.012 0 0 0
60 0.980 0.020 0 0 0
70 0.994 0.006 0 0 0
80 0.996 0.002 0.002 0 0
90 0.996 0.002 0.002 0 0
100 0.996 0.004 0 0 0
110 0.994 0.006 0 0 0
120 0.990 0.010 0 0 0
130 0.994 0.006 0 0 0
140 0.994 0.006 0 0 0
150 0.994 0.006 0 0 0
Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment I
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.004 0 0 0.484 0.512
20 0 0 0 0.326 0.674
30 0 0 0 0.194 0.806
40 0 0 0 0.090 0.910
50 0 0 0 0.076 0.924
60 0 0 0 0.030 0.970
70 0 0 0 0.012 0.988
80 0 0 0 0.010 0.990
90 0 0 0 0.006 0.994
100 0 0 0 0.004 0.996
110 0 0 0 0.006 0.994
120 0 0 0 0 1
130 0 0 0 0 1
140 0 0 0 0.002 0.998
150 0 0 0 0 1
Figure 1: The sample means and error bars of $\hat{K}$ in Experiment I
Table 5: Empirical rejection rates in Experiment II, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
$H_{0}$ $H_{1}$
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
10 0.042 0.050 0.046 0.042 0.052 0.076 0.048 0.046 0.998 0.996 0.998 0.996 0.998 0.998
20 0.034 0.042 0.026 0.050 0.060 0.070 0.054 0.018 1 1 1 1 1 1
30 0.028 0.048 0.036 0.046 0.046 0.070 0.062 0.022 1 1 1 1 1 1
40 0.054 0.054 0.046 0.050 0.060 0.062 0.052 0.042 1 1 1 1 1 1
50 0.064 0.062 0.042 0.060 0.050 0.068 0.058 0.032 1 1 1 1 1 1
60 0.032 0.054 0.060 0.040 0.040 0.078 0.064 0.026 1 1 1 1 1 1
70 0.076 0.038 0.052 0.052 0.036 0.052 0.042 0.022 1 1 1 1 1 1
80 0.040 0.046 0.034 0.048 0.058 0.074 0.068 0.028 1 1 1 1 1 1
90 0.044 0.052 0.050 0.054 0.042 0.060 0.050 0.032 1 1 1 1 1 1
100 0.038 0.068 0.050 0.046 0.046 0.064 0.056 0.028 1 1 1 1 1 1
110 0.052 0.052 0.046 0.052 0.048 0.050 0.046 0.032 1 1 1 1 1 1
120 0.052 0.042 0.038 0.034 0.052 0.040 0.036 0.030 1 1 1 1 1 1
130 0.048 0.052 0.044 0.036 0.048 0.062 0.054 0.040 1 1 1 1 1 1
140 0.062 0.058 0.056 0.038 0.038 0.046 0.042 0.044 1 1 1 1 1 1
150 0.038 0.056 0.046 0.052 0.060 0.060 0.054 0.034 1 1 1 1 1 1
Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment II
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.968 0.026 0.004 0 0.002
20 0.982 0.016 0.002 0 0
30 0.978 0.020 0.002 0 0
40 0.992 0.008 0 0 0
50 0.990 0.006 0.004 0 0
60 0.996 0.004 0 0 0
70 0.994 0.006 0 0 0
80 0.984 0.016 0 0 0
90 0.996 0.004 0 0 0
100 0.994 0.006 0 0 0
110 0.996 0.004 0 0 0
120 0.996 0.004 0 0 0
130 0.994 0.006 0 0 0
140 0.994 0.006 0 0 0
150 0.988 0.012 0 0 0
Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment II
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.002 0.984 0.012 0 0.002
20 0 0.994 0.004 0 0.002
30 0 0.990 0.008 0 0.002
40 0 0.996 0.004 0 0
50 0 0.996 0.004 0 0
60 0 0.994 0.004 0 0.002
70 0 0.998 0.002 0 0
80 0 1 0 0 0
90 0 1 0 0 0
100 0 0.996 0.004 0 0
110 0 0.998 0.002 0 0
120 0 0.994 0.006 0 0
130 0 0.998 0.002 0 0
140 0 0.998 0.002 0 0
150 0 0.996 0.004 0 0
Figure 2: The sample means and error bars of $\hat{K}$ in Experiment II
The next two experiments are designed to evaluate the tests for models (3.9) and (3.10), respectively. All settings are inherited from the previous two experiments, except that the data-generating process is modified accordingly.
@@
Here $\{Y_{ij}^{(0,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$. Note that in Experiment III, $\mathrm{E}[Y_{ij}^{(0,3)}]=\mathrm{E}[Y_{ij}^{(1,3)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,3)}]=\operatorname{Var}[Y_{ij}^{(1,3)}]=4$, which corresponds to model (3.9); in Experiment IV, $\mathrm{E}[Y_{ij}^{(0,4)}]=\mathrm{E}[Y_{ij}^{(1,4)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,4)}]=\operatorname{Var}[Y_{ij}^{(1,4)}]=j^{2}$, which corresponds to model (3.10). The conclusions from Experiments III and IV are analogous to those from Experiments I and II, respectively (see Tables 813 and Figures 34), thus validating our unified testing framework. To examine the robustness of our findings with respect to the number of groups, we additionally consider a variant of Experiment III with $J=10$ (referred to as Experiment III). The results corroborate the main conclusions, as reported in A.
Table 8: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
$10$ 0.050 0.046 0.048 0.048 0.052 0.076 0.050
2020 0.072 0.060 0.050 0.046 0.050 0.054 0.048
$30$ 0.054 0.046 0.054 0.038 0.056 0.074 0.062
$40$ 0.046 0.048 0.060 0.054 0.048 0.054 0.050
$50$ 0.044 0.060 0.052 0.052 0.048 0.040 0.038
$60$ 0.066 0.044 0.048 0.052 0.058 0.048 0.042
$70$ 0.044 0.048 0.038 0.046 0.048 0.072 0.064
$80$ 0.050 0.044 0.048 0.038 0.042 0.074 0.068
$90$ 0.040 0.060 0.046 0.054 0.046 0.048 0.042
$100$ 0.062 0.056 0.042 0.050 0.048 0.058 0.052
$110$ 0.060 0.046 0.066 0.052 0.050 0.068 0.054
$120$ 0.034 0.044 0.038 0.038 0.036 0.058 0.056
$130$ 0.038 0.050 0.042 0.052 0.050 0.058 0.048
$140$ 0.038 0.050 0.048 0.042 0.044 0.044 0.040
$150$ 0.064 0.056 0.048 0.058 0.058 0.046 0.042
Table 9: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.976 0.024 0 0 0
2020 0.974 0.018 0.006 0 0.002
$30$ 0.986 0.014 0 0 0
$40$ 0.990 0.010 0 0 0
$50$ 0.998 0.002 0 0 0
$60$ 0.988 0.012 0 0 0
$70$ 0.992 0.008 0 0 0
$80$ 0.990 0.008 0.002 0 0
$90$ 0.988 0.012 0 0 0
$100$ 0.996 0.004 0 0 0
$110$ 0.994 0.004 0.002 0 0
$120$ 0.992 0.006 0.002 0 0
$130$ 0.998 0.002 0 0 0
$140$ 0.992 0.008 0 0 0
$150$ 0.996 0.004 0 0 0
Table 10: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.032 0.004 0 0.418 0.546
2020 0 0 0 0.314 0.686
$30$ 0 0 0 0.174 0.826
$40$ 0 0 0 0.098 0.902
$50$ 0 0 0 0.064 0.936
$60$ 0 0 0 0.036 0.964
$70$ 0 0 0 0.022 0.978
$80$ 0 0 0 0.018 0.982
$90$ 0 0 0 0.008 0.992
$100$ 0 0 0 0.008 0.992
$110$ 0 0 0 0.002 0.998
$120$ 0 0 0 0 1
$130$ 0 0 0 0 1
$140$ 0 0 0 0.002 0.998
$150$ 0 0 0 0 1
Figure 3: The sample means and error bars of $\hat{K}$ in Experiment III
Figure 4: The sample means and error bars of $\hat{K}$ in Experiment IV

### v4

@@ 5 Simulations
The results of Experiment II are quite similar; see Tables 4, 5, 6, and Figure 2. However, from Table 6, under $H_{1}$, the test statistic $N\widehat{\Psi}_{1}^{2}$ fails to exhibit power as a result of $\mathbb{E}[\pi_{1}(Z)]=\mathbb{E}[\pi_{1}(\Phi(Y_{ij}^{(1,2)}-8))]=0$ in this case. This illustrates that $N\widehat{\Psi}_{K}^{2}$ may be inconsistent against certain alternatives, especially when $K$ is small. Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, $\{\widehat{K}=1\}$ does not occur, consistent with Theorem 10 (1), which ensures that $\mathbb{P}(\widehat{K}\geq K_{0})\to 1$ under $H_{1}^{\prime}$ with $K_{0}=2$.
Table 2: Empirical rejection rates under $H_{0}$ in Experiment I. The empirical power under $H_{1}$ equals 1 for all sample sizes and all considered tests.
Smooth tests SW JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.050 0.048 0.048 0.044 0.040 0.056 0.034 0.054 0.044 0
20 0.046 0.044 0.034 0.040 0.048 0.066 0.046 0.054 0.042 0
30 0.042 0.036 0.044 0.068 0.068 0.060 0.046 0.040 0.032 0.002
40 0.058 0.040 0.036 0.058 0.052 0.058 0.050 0.044 0.038 0
50 0.064 0.044 0.054 0.054 0.062 0.068 0.060 0.036 0.030 0
60 0.054 0.034 0.038 0.044 0.050 0.070 0.062 0.056 0.058 0
70 0.056 0.038 0.048 0.046 0.052 0.056 0.048 0.072 0.062 0
80 0.038 0.050 0.072 0.042 0.048 0.058 0.048 0.046 0.038 0
90 0.062 0.066 0.052 0.064 0.058 0.048 0.040 0.028 0.038 0
100 0.034 0.050 0.050 0.060 0.054 0.052 0.038 0.038 0.040 0
110 0.054 0.032 0.054 0.060 0.050 0.038 0.036 0.042 0.044 0
120 0.052 0.052 0.038 0.044 0.052 0.052 0.050 0.050 0.046 0
130 0.052 0.050 0.064 0.060 0.058 0.068 0.064 0.050 0.060 0
140 0.062 0.040 0.026 0.046 0.030 0.038 0.030 0.068 0.050 0
150 0.054 0.044 0.062 0.034 0.058 0.056 0.048 0.064 0.066 0
Table 3: Empirical frequency of $\widehat{K}$ in Experiment I
Under $H_{0}$ Under $H_{1}$
$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
10 0.982 0.018 0 0 0 0.004 0 0 0.484 0.512
20 0.988 0.010 0 0 0.002 0 0 0 0.326 0.674
30 0.986 0.012 0 0.002 0 0 0 0 0.194 0.806
40 0.984 0.014 0.002 0 0 0 0 0 0.090 0.910
50 0.988 0.012 0 0 0 0 0 0 0.076 0.924
60 0.980 0.020 0 0 0 0 0 0 0.030 0.970
70 0.994 0.006 0 0 0 0 0 0 0.012 0.988
80 0.996 0.002 0.002 0 0 0 0 0 0.010 0.990
90 0.996 0.002 0.002 0 0 0 0 0 0.006 0.994
100 0.996 0.004 0 0 0 0 0 0 0.004 0.996
110 0.994 0.006 0 0 0 0 0 0 0.006 0.994
120 0.990 0.010 0 0 0 0 0 0 0 1
130 0.994 0.006 0 0 0 0 0 0 0 1
140 0.994 0.006 0 0 0 0 0 0 0.002 0.998
150 0.994 0.006 0 0 0 0 0 0 0 1
Figure 1: The sample means and error bars of $\widehat{K}$ in Experiment I
Table 4: Empirical rejection rates under $H_{0}$ in Experiment II
Smooth tests SW JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.042 0.050 0.046 0.042 0.052 0.076 0.048 0.050 0.042 0
20 0.034 0.042 0.026 0.050 0.060 0.070 0.054 0.032 0.042 0
30 0.028 0.048 0.036 0.046 0.046 0.070 0.062 0.064 0.052 0.002
40 0.054 0.054 0.046 0.050 0.060 0.062 0.052 0.044 0.048 0
50 0.064 0.062 0.042 0.060 0.050 0.068 0.058 0.042 0.044 0
60 0.032 0.054 0.060 0.040 0.040 0.078 0.064 0.056 0.062 0
70 0.076 0.038 0.052 0.052 0.036 0.052 0.042 0.054 0.054 0
80 0.040 0.046 0.034 0.048 0.058 0.074 0.068 0.044 0.044 0
90 0.044 0.052 0.050 0.054 0.042 0.060 0.050 0.052 0.052 0
100 0.038 0.068 0.050 0.046 0.046 0.064 0.056 0.044 0.038 0
110 0.052 0.052 0.046 0.052 0.048 0.050 0.046 0.046 0.060 0
120 0.052 0.042 0.038 0.034 0.052 0.040 0.036 0.044 0.052 0
130 0.048 0.052 0.044 0.036 0.048 0.062 0.054 0.054 0.038 0
140 0.062 0.058 0.056 0.038 0.038 0.046 0.042 0.032 0.044 0
150 0.038 0.056 0.046 0.052 0.060 0.060 0.054 0.050 0.050 0
Table 5: Empirical rejection rates under $H_{1}$ in Experiment II
Smooth tests SW JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.046 0.998 0.996 0.998 0.996 0.998 0.998 1 0.990 0.824
20 0.018 1 1 1 1 1 1 1 1 0.994
30 0.022 1 1 1 1 1 1 1 1 1
40 0.042 1 1 1 1 1 1 1 1 1
50 0.032 1 1 1 1 1 1 1 1 1
60 0.026 1 1 1 1 1 1 1 1 1
70 0.022 1 1 1 1 1 1 1 1 1
80 0.028 1 1 1 1 1 1 1 1 1
90 0.032 1 1 1 1 1 1 1 1 1
100 0.028 1 1 1 1 1 1 1 1 1
110 0.032 1 1 1 1 1 1 1 1 1
120 0.030 1 1 1 1 1 1 1 1 1
130 0.040 1 1 1 1 1 1 1 1 1
140 0.044 1 1 1 1 1 1 1 1 1
150 0.034 1 1 1 1 1 1 1 1 1
Table 6: Empirical frequency of $\widehat{K}$ in Experiment II
Under $H_{0}$ Under $H_{1}$
$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
10 0.968 0.026 0.004 0 0.002 0.002 0.984 0.012 0 0.002
20 0.982 0.016 0.002 0 0 0 0.994 0.004 0 0.002
30 0.978 0.020 0.002 0 0 0 0.990 0.008 0 0.002
40 0.992 0.008 0 0 0 0 0.996 0.004 0 0
50 0.990 0.006 0.004 0 0 0 0.996 0.004 0 0
60 0.996 0.004 0 0 0 0 0.994 0.004 0 0.002
70 0.994 0.006 0 0 0 0 0.998 0.002 0 0
80 0.984 0.016 0 0 0 0 1 0 0 0
90 0.996 0.004 0 0 0 0 1 0 0 0
100 0.994 0.006 0 0 0 0 0.996 0.004 0 0
110 0.996 0.004 0 0 0 0 0.998 0.002 0 0
120 0.996 0.004 0 0 0 0 0.994 0.006 0 0
130 0.994 0.006 0 0 0 0 0.998 0.002 0 0
140 0.994 0.006 0 0 0 0 0.998 0.002 0 0
150 0.988 0.012 0 0 0 0 0.996 0.004 0 0
Figure 2: The sample means and error bars of $\widehat{K}$ in Experiment II
The next two experiments, focusing on models (3.9) and (3.10), follow simulation settings similar to the previous ones. The data-generating process for each model is specified as follows.
@@
Here $\{Y_{ij}^{(0,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(0,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$, while $\{Y_{ij}^{(1,3)}\}_{i=1,j=1}^{jm,5}$ and $\{Y_{ij}^{(1,4)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{1}$. Note that in Experiment III, $\mathbb{E}[Y_{ij}^{(0,3)}]=\mathbb{E}[Y_{ij}^{(1,3)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,3)}]=\operatorname{Var}[Y_{ij}^{(1,3)}]=4$, which corresponds to model (3.9); in Experiment IV, $\mathbb{E}[Y_{ij}^{(0,4)}]=\mathbb{E}[Y_{ij}^{(1,4)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,4)}]=\operatorname{Var}[Y_{ij}^{(1,4)}]=j^{2}$, which corresponds to model (3.10). For the smooth tests, the results of Experiments III and IV are broadly consistent with those of Experiments I and II (see Tables 711 and Figures 34), further supporting the validity of our unified testing framework. For the other three classical methods, a notable phenomenon arises in Experiment IV: under $H_{0}$, the Jarque–Bera test exhibits undersizing, suggesting that it is not well suited for models with heterogeneous group variances. In contrast, our proposed test remains well-calibrated and maintains stable size control.